OpenStudy (anonymous):

A number is divided by 2. If 16 is added to the quotient you get 20.

OpenStudy (anonymous):

4 :)

OpenStudy (anonymous):

I mean 8,, not four sorry :(

OpenStudy (anonymous):

4/2=2+16=18 ??

OpenStudy (anonymous):

yeah 8

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

A number is multiplied by 3. If 3 is subtracted from the product you get 24.

OpenStudy (anonymous):

9

OpenStudy (anonymous):

A number is multiplied by 6. If 8 is subtracted from the product you get 34.

OpenStudy (anonymous):

how did yall get 8

OpenStudy (texaschic101):

x/2 + 16 = 20 (multiply the equation by 2) x + 32 = 40 (subtract 32 from each side) x = 40 - 32 x = 8

OpenStudy (anonymous):

A number is multiplied by 6. If 8 is subtracted from the product you get 34.

OpenStudy (texaschic101):

a number is multiplied by 6, if 8 is subtracted from the product, you get 34. 6x - 8 = 34 (add 8 to each side) 6x = 34 + 8 6x = 42 x = 7

OpenStudy (anonymous):

A number is divided by 3. If 8 is subtracted from the quotient you get -6.

OpenStudy (texaschic101):

x/3 - 8 = -6 (multiply the equation by 3) x - 24 = - 18 (add 24 to each side) x = -18 + 24 x = 6

OpenStudy (anonymous):

What is the slope of the line that contains the points (2, 6) and (-1, -6)?

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) (2,6) x1 = 2 and y1 = 6 (-1,-6) x2 = -1 and y2 = -6 now lets sub slope(m) = (-6 - 6) / (-1 - 2) m = -12/-3 m = 4 <== slope is 4

OpenStudy (anonymous):

What is the slope of the line that contains the points (-2, 1) and (0, -3)?

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) m = (-3 - 1) / (0 - (-2) m = -4 / 2 m = -2 <== slope is -2

OpenStudy (anonymous):

What is the slope of the line that contains the points (-6, -7) and (6, -5)?

OpenStudy (anonymous):

1/11

OpenStudy (texaschic101):

slope(m) = )y2 - y1) / (x2 - x1) m = (-5 - (-7) / (6 - (-6) m = (-5 + 7) / (6 + 6) m = 2/12 m = 1/6 <== slope is 1/6

OpenStudy (anonymous):

Gmexi_Girl YOU WRONG texaschic101 you right

OpenStudy (texaschic101):

be nice...she tried. Did you ?

OpenStudy (anonymous):

I was being nice and yes I did tried BIG HOE

OpenStudy (anonymous):

.....

OpenStudy (texaschic101):

I am done here. Your on you own